Please help me. I'm struggling with these word problems. This is too difficult. Diane has $1.65 in dimes and nickels. She has a total of 21 coins. How many of each does she have?
Suppose Tom has 14 coins totaling $2.30. If he has on dimes and quarters, how many of each does he have?

Answer by Edwin McCravy(20086)

Start with the easy sentence first:

"She has a total of 21 coins."
Let d = the number of dime coins and n = the number of nickel coins.
Then if you add them together you put a + sign between them like this

d+n

and since the total number of coins is 21, we have the equation

d+n = 21

Now look at the other clause:

"...$1.65 in dimes and nickels."
If you multiply the number of dimes, d, by 10 cents you get 10d cents
If you multiply the number of nickels, n, by 5 cents you get 5n cents
If you add 10d cents and 6n cents, you put a + sign between them, like
this:

10d+5n

That expression must represent the total number of cents. 
But you also know that the total number of cents in $1.65 is 165 cents.
So 10d+5n must equal to 165.  So you have the equation

    10d+5n = 165

So you have the systme of equations

d+n = 21
10d+5n = 165

which you can solve by substitution or elimination and get d=12, n=9

If you don't know how to solve that system of equations, then
post again asking how.

----------------------

Suppose Tom has 14 coins totaling $2.30. If he has only dimes and quarters,
how many of each does he have?
The exact same way. You'll get d=8, q=6

Edwin